a certain recipe asks for 3/5 cup of granola. how much is needed to make 3/4 of a recipe?

Answer by TimothyLamb(4379)

---
(3/5)(3/4) = 9/20 = 0.45 cup of granola
---
note that a full recipe asks for: 3/5 = 0.6 cup of granola
---
and note that:
(0.6)(3/4) = (1.8)/4 = 0.45 cup
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php